# Simulating the Visual World with Artificial Intelligence: A Roadmap

### v2

@@ Abstract
The landscape of video generation is shifting, from a focus on generating visually appealing clips to building virtual environments that support interaction and maintain physical plausibility. These developments point toward the emergence of video foundation models that function not only as visual generators but also as implicit world models, models that simulate the physical dynamics, agent-environment interactions, and task planning that govern real or imagined worlds. This survey provides a systematic overview of this evolution, conceptualizing modern video foundation models as the combination of two core components: an implicit world model and a video renderer. The world model encodes structured knowledge about the world, including physical laws, interaction dynamics, and agent behavior. It serves as a latent simulation engine that enables coherent visual reasoning, long-term temporal consistency, and goal-driven planning. The video renderer transforms this latent simulation into realistic visual observations, effectively producing videos as a “window” into the simulated world. We trace the progression of video generation through four generations, in which the core capabilities advance step by step, ultimately culminating in a world model, built upon a video generation model, that embodies intrinsic physical plausibility, real-time multimodal interaction, and planning capabilities spanning multiple spatiotemporal scales. For each generation, we define its core characteristics, highlight representative works, and examine their application domains such as robotics, autonomous driving, and interactive gaming. Finally, we discuss open challenges and design principles for next-generation world models, including the role of agent intelligence in shaping and evaluating these systems.
@@ 1.1 Motivation
World models, which aim to simulate the real world, have long posed significant challenges in artificial intelligence, influencing various applications, such as robotics, autonomous driving, and gaming. While the specific capabilities required for world models are numerous and not yet precisely defined, approaches such as 3D generation [lumagenie2025, liu2025unleashing, huang2024placiddreamer, yang2019pointflow, cao2025physx], 3D/4D scene generation [sun20253d, hsu2025programs, yang2024holodeck, yu2025wonderworld, yu2024wonderjourney, huang2025omnix, huang2025terra, zhaoxi20254dnex], and video generation demonstrate one or more relevant abilities, such as motion dynamics, interaction and controllability, visual quality, 3D consistency, and generation efficiency. Consequently, these approaches have been adopted in recent works as potential pathways toward realizing world models.
While these approaches each excel in partial capabilities toward world modeling, video generation offers one direct and comprehensive pathway, which may be a promising tool for building a world model. From a cognitive science perspective, vision is the dominant sensory modality through which both humans and embodied agents perceive, learn, and reason about the world. Visual streams not only convey spatial layout and object properties but also encode temporal dynamics and causal relationships crucial for prediction and planning. Even complex 3D or 4D simulations can be rendered into videos or images for interpretation, meaning that any human or embodied agent grounds its understanding in visual sequences. This intrinsic reliance on visual representation makes video generation a uniquely natural and information-rich foundation for constructing world models.
Recent advances in end-to-end video generation indicate that such models can now serve as high-quality visual renderers, enabling the exploration of video-based approaches to world modeling. Recent advancements in techniques, such as diffusion models [peebles2023dit, podell2023sdxl, esser2024sd3, song2020denoising] and autoregressive transformers [parmar2018imagegeneration, ramesh2021zeroshott2i, razavi2019vqvae2], have made it possible to generate high-quality videos grounded in fundamental world knowledge. Current methods [sora, teng2025magi, wan2025wan, peng2025opensora2, kong2024hunyuanvideo, xu2024easyanimate, bao2024vidu, yin2025causvid, polyak2024moviegen, huang2025selfforcing, klingai2024kling, yang2024cogvideox, lin2025apt2, agarwal2025cosmos, chen2025skyreels, gao2025seedance, mirage2025] based on these backbones are now capable of producing long-duration, high-quality videos with superficial faithfulness, endowing them with the ability to simulate real-world environments with high fidelity and incorporate multi-modal conditioning. As a result, there is an increasingly evident trend of utilizing video generation models as world models.
@@
In this survey, we define a physical world model as a sophisticated digital engine that encodes comprehensive world knowledge to simulate real‑world dynamics in accordance with intrinsic physical and mathematical laws. Such models serve both as high‑fidelity simulators for advancing domains like robotics, autonomous driving, and gaming, and as controlled testbeds for training and evaluating intelligent agents under realistic yet safe conditions. By modeling the physical world, they can also support physics-based engineering, high-stakes decision-making, and other real-world tasks.
@@
Despite these advances, challenges remain at both the conceptual and structural levels. Conceptually, the definition of a “world model” remains ambiguous, making it difficult to unify perspectives and evaluate progress. Structurally, the field lacks a well-established taxonomy to organize modeling capabilities, developmental stages, and potential trajectories. These gaps highlight an urgent need for a systematic elucidation and comprehensive survey to consolidate existing knowledge and guide the next stage of research. Current surveys and benchmarks [kong20253d, wang2025survey, xing2024survey, yu2025surveyigv, dal2024jointadsurvey, zhu2024issoraasurvey, fu2024exploringsurvey, melnik2024videosurvey, lin2025exploringsurvey, guan2024adinitialsurvey, liu2025generativesurvey, cho2024soraasansgisurvey, sun2024fromsorawecanseesurvey, yu2025position, duan2025worldscore, qin2024worldsimbench, li2025worldmodelbench, zheng2025vbench2, zhang2025evaluationagent] have laid important groundwork by summarizing related methodologies, datasets, and applications. However, there remains a need to explicitly clarify what aspects have been thoroughly addressed, and which areas, such as real-time integration, controllable video-to-world pipelines, or comprehensive evaluation metrics, are still underexplored. Accordingly, this survey aims to chart a clear path from video generation toward comprehensive world modeling, providing guidance for future research and development in this emerging field.
In this paper, we begin with formal definitions and provide a detailed discussion centered around the taxonomy of world models. We define a physical world model as a digital simulation engine embedded with comprehensive world knowledge, capable of predicting the next scene conditioned on environmental states and contextual priors. Each prediction step can be represented as a triplet (Current Scene,Navigation Mode,Prior Information), which collectively determine the evolution of the simulated environment. The model focuses on capturing the causal and spatiotemporal dynamics of the physical environment, while external inputs such as navigation modes or actions act as stimuli that perturb or impact the environment’s evolution. Hence, the world model is formulated as an interactive environment system that responds to external interventions without explicitly modeling the decision-making process that generates them. Based on this definition, we systematically analyze how video generation models have evolved toward world modeling and propose a four-generation taxonomy according to model capability, as illustrated in Figure 1.
@@ 3 Roadmap
Semantic and Navigational Interactiveness. Interactiveness is the core capability emphasized in Generation 2, termed as semantic and navigational interactiveness, marking a significant advancement in the dimensions of control flexibility, condition-video consistency, subject-centric controllability, and intrinsic model competence. Models in this generation enable flexible control, particularly conditioned on navigation modes, including actions, text commands, and predefined trajectories. They also exhibit emerging basic reasoning abilities, such as inferring motion sequences from high-level commands. In terms of condition-video consistency, Generation 2 models demonstrate notable improvements: they can generate complete and coherent motions with minimal dynamic distortion, and maintain visually appropriate and semantically aligned backgrounds. Another hallmark of this generation is basic subject-centric external control, where the model is capable of interpreting and executing control signals directed at a specific subject, such as instructing one agent to perform a sequence of actions or dynamically adjusting the viewpoint around that subject. However, these models typically operate within relatively static or simplified backgrounds. In addition, the intrinsic model capabilities have advanced significantly, as evidenced by the integration of 3D spatial understanding and stronger semantic comprehension rather than basic pixel-level understanding. Together, these advancements position Generation 2 as a transitional phase toward truly interactive and semantically grounded world modeling.
@@
Furthermore, Generation 4 achieves arbitrary spatial and temporal scale planning. In the spatial domain, the model can plan across macroscopic scales such as universe-level evolution and microscopic scales like microbial dynamics or atomic-level transitions. Similarly, in the temporal domain, the model is capable of operating across vast time scales, from long-term evolution spanning years or centuries (requiring time compression and critical event selection ability), to mid-scale physical world dynamics, down to fine-grained, high-frequency phenomena such as insect wing beats or human pupil micro-movements. This ability to plan across stochastic events and arbitrary scales represents a critical step toward building general-purpose simulation engines that align more closely with the complexity and uncertainty of the real world.
@@ 4 Methods: A Hierarchical Taxonomy
Before delving into each generation in detail, it is important to emphasize that the three core capabilities, faithfulness, interactiveness, and planning, develop in parallel rather than appearing sequentially. Each generation highlights one capability as its dominant focus, reflecting the primary research emphasis and maturity level of that stage, yet all three capabilities continue to advance simultaneously across generations. This co-evolution reflects the natural progression of video generation models toward comprehensive world modeling, where visual realism, controllability, and long-horizon reasoning are continuously refined in tandem rather than introduced one after another.
Similarly, the application domains, general scenes, robotics, autonomous driving, and gaming, analyzed in this section are not exclusive to any single generation. Instead, they overlap across all four, as the underlying modeling principles remain consistent while the level of capability and control gradually increases. We organize these applications under Generation 2 primarily because most of the recent works and open-source systems lie in this stage. Subsequent generations naturally extend these same applications toward higher-level planning, real-time interaction, and stochastic world simulation.
@@ 4.1.1 Foundation Models with Visual-Centric World Knowledge
We summarize representative pre-trained text-to-video (T2V) models [hong2022cogvideo, chen2023videocrafter1, chen2024videocrafter2, blattmann2023svd, wang2023modelscope, gen223, guo2023animatediff, zhang2024show1, wang2025lavie, qing2024higen, ma2024latte, wang2024magicvideov2, bar2024lumiere, klingai2024kling, henschel2025streamingt2v, girdhar2023emuvideo, kondratyuk2023videopoet, yuan2024instructvideo, zhou2024allegro, tian2024videotetris], along with commonly used pre-trained image-to-video (I2V) models [zhang2023i2vgen, lin2024opensoraplan, hailuo24minmax, xu2024easyanimate, klingai2024kling, blattmann2023svd, yu2023magvitv2, chen2023videocrafter1, sun2024dimensionx, kondratyuk2023videopoet, guo2024i2vadapter, team2025kling], examining their backbone architectures, video lengths, resolutions, and post-training capabilities related to interactivity and embodiment which are core abilities for serving as world models. This survey also covers both open-sourced models [ma2024latte, wang2025lavie, xu2024easyanimate, si2025repvideo, chen2024videocrafter2, guo2023animatediff, hacohen2024ltxvideo, chen2023videocrafter1, zheng2024opensora, zhang2024show1, lin2024opensoraplan, qing2024higen, wang2023modelscope, henschel2025streamingt2v, sun2024dimensionx, kondratyuk2023videopoet, zhou2024allegro] and commercial APIs [hailuo24minmax, klingai2024kling, gen223, pika23, girdhar2023emuvideo]. For instance, CogVideo [hong2022cogvideo] is widely regarded as an early open-source, large-scale pretrained T2V model that played a pivotal role in initiating the development of video foundation models. The release of Kling 1.0 [klingai2024kling], built on a DiT backbone, further advanced generative AI toward longer duration and greater visual quality.
Following the initial success of Generation 1 models in producing basic text–video consistency and plausible motion dynamics, some research shifted focus toward long and efficient video generation, sometimes at the expense of faithfulness. This led to the development of streaming video generation based on an autoregressive architecture. Specifically, LTX‑Video [hacohen2024ltxvideo] achieves semi‑real‑time generation, continuously producing video clips faster than they can be viewed, by leveraging a high‑compression Video‑VAE and a denoising transformer with full spatiotemporal attention. While notable for its generation speed, its limited fidelity of generated videos still places it within the first generation.
@@ 4.1.3 Navigation World Model
Navigation world models inviting $N$ into the input set as $I=\{T,O,Au,N\}$. In general scenes, navigation world models are primarily developed for motion control tasks, including local subject motion [burgert2025gowiththeflow, geng2025motionprompting, namekata2024sgi2v, ma2024trailblazer, jain2024peekaboo, wang2024motionctrl, wu2024draganything, qiu2024freetraj, yang2024directavideo] and camera motion [bai2025recammaster, wang2025cinemaster, bahmani2025ac3d, bahmani2024vd3d, he2024cameractrl, van2024gcd, kuang2024cvd, bai2024syncammaster]. These methods often build upon pre-trained video foundation models [blattmann2023svd, wang2023modelscope, chen2023videocrafter1, yang2024cogvideox, bar2024lumiere, guo2023animatediff], incorporating training-free control modules to enable interactive motion and camera control. However, due to current limitations in video length, generation quality, and the diversity of controllable entities, these approaches remain in the Generation 1 world models. For local motion control tasks, DragAnything [wu2024draganything] offers intuitive point‑based dragging to finely adjust subject motion, enabling precise spatial manipulation in generated videos. Peekaboo [jain2024peekaboo] elevates this to semantic‑level control, where object‑centric motion can be directed in line with textual prompts. Pushing further, Trailblazer [ma2024trailblazer] incorporates trajectory‑driven navigation, supporting complex and continuous motion generation for both subjects and cameras, advancing toward more coherent and context‑aware video navigation. In terms of camera control, existing methods [bahmani2024vd3d, bahmani2025ac3d, xu2024camco] inject camera motion trajectories, represented in Plücker coordinates, into pre-trained foundation models via ControlNet-based modules. Alternatively, methods such as GCD [van2024gcd] directly manipulate camera extrinsics, i.e., rotation and translation, as input to control camera viewpoints.

### v3

@@ Abstract
The landscape of video generation is shifting, from a focus on generating visually appealing clips to building virtual environments that support interaction and maintain physical plausibility. These developments point towards the emergence of video foundation models that function not only as visual generators but also as implicit world models, models that simulate the physical dynamics, agent-environment interactions, and task planning that govern real or imagined worlds. This survey provides a systematic overview of this evolution, conceptualizing modern video foundation models as the combination of two core components: an implicit world model and a video renderer. The world model encodes structured knowledge about the world, including physical laws, interaction dynamics, and agent behavior. It serves as a latent simulation engine that enables coherent visual reasoning, long-term temporal consistency, and goal-driven planning. The video renderer transforms this latent simulation into realistic visual observations, effectively producing videos as a “window” into the simulated world. We trace the progression of video generation through four generations, in which the core capabilities advance step by step, ultimately culminating in a world model, built upon a video generation model, that embodies intrinsic physical plausibility, real-time multimodal interaction, and planning capabilities spanning multiple spatiotemporal scales. For each generation, we define its core characteristics, highlight representative works, and examine their application domains such as robotics, autonomous driving, and interactive gaming. Finally, we discuss open challenges and design principles for next-generation world models, including the role of agent intelligence in shaping and evaluating these systems.
@@ 1.1 Motivation
World models, which aim to simulate the real world, have long posed significant challenges in artificial intelligence, influencing various applications, such as robotics, autonomous driving, and gaming. While the specific capabilities required for world models are numerous and not yet precisely defined, approaches such as 3D generation [lumagenie2025, liu2025unleashing, huang2024placiddreamer, yang2019pointflow, cao2025physx], 3D/4D scene generation [sun20253d, hsu2025programs, yang2024holodeck, yu2025wonderworld, yu2024wonderjourney, huang2025omnix, huang2025terra, zhaoxi20254dnex], and video generation demonstrate one or more relevant abilities, such as motion dynamics, interaction and controllability, visual quality, 3D consistency, and generation efficiency. Consequently, these approaches have been adopted in recent works as potential pathways towards realizing world models.
While these approaches each excel in partial capabilities towards world modeling, video generation offers one direct and comprehensive pathway, which may be a promising tool for building a world model. From a cognitive science perspective, vision is the dominant sensory modality through which both humans and embodied agents perceive, learn, and reason about the world. Visual streams not only convey spatial layout and object properties but also encode temporal dynamics and causal relationships crucial for prediction and planning. Even complex 3D or 4D simulations can be rendered into videos or images for interpretation, meaning that any human or embodied agent grounds its understanding in visual sequences. This intrinsic reliance on visual representation makes video generation a uniquely natural and information-rich foundation for constructing world models.
Recent advances in end-to-end video generation indicate that such models can now serve as high-quality visual renderers, enabling the exploration of video-based approaches to world modeling. Recent advancements in techniques, such as diffusion models [peebles2023dit, podell2023sdxl, esser2024sd3, song2020denoising] and autoregressive transformers [parmar2018imagegeneration, razavi2019vqvae2], have made it possible to generate high-quality videos grounded in fundamental world knowledge. Current methods [sora, teng2025magi, wan2025wan, peng2025opensora2, kong2024hunyuanvideo, xu2024easyanimate, bao2024vidu, yin2025causvid, polyak2024moviegen, huang2025selfforcing, klingai2024kling, yang2024cogvideox, lin2025apt2, agarwal2025cosmos, chen2025skyreels, gao2025seedance, mirage2025] based on these backbones are now capable of producing long-duration, high-quality videos with superficial faithfulness, endowing them with the ability to simulate real-world environments with high fidelity and incorporate multi-modal conditioning. As a result, there is an increasingly evident trend of utilizing video generation models as world models.
@@
In this survey, we define a physical world model as a digital engine that encodes comprehensive world knowledge to simulate real‑world dynamics according to intrinsic physical laws. Such models serve both as high‑fidelity simulators for advancing domains like robotics, autonomous driving, and gaming, and as controlled testbeds for training and evaluating intelligent agents under realistic conditions.
@@
Despite these advances, challenges remain at both the conceptual and structural levels. Conceptually, the definition of a “world model” remains ambiguous, making it difficult to unify perspectives and evaluate progress. Structurally, the field lacks a well-established taxonomy to organize modeling capabilities, developmental stages, and potential trajectories. These gaps highlight an urgent need for a systematic elucidation and comprehensive survey to consolidate existing knowledge and guide the next stage of research. Current surveys and benchmarks [kong20253d, wang2025survey, xing2024survey, yu2025surveyigv, dal2024jointadsurvey, zhu2024issoraasurvey, fu2024exploringsurvey, melnik2024videosurvey, lin2025exploringsurvey, guan2024adinitialsurvey, liu2025generativesurvey, cho2024soraasansgisurvey, sun2024fromsorawecanseesurvey, yu2025position, duan2025worldscore, qin2024worldsimbench, li2025worldmodelbench, zheng2025vbench2, zhang2025evaluationagent] have laid important groundwork by summarizing related methodologies, datasets, and applications. However, there remains a need to explicitly clarify what aspects have been thoroughly addressed, and which areas, such as real-time integration, controllable video-to-world pipelines, or comprehensive evaluation metrics, are still underexplored. Accordingly, this survey aims to chart a clear path from video generation towards comprehensive world modeling, providing guidance for future research and development in this emerging field.
In this paper, we begin with formal definitions and provide a detailed discussion centered around the taxonomy of world models. We define a physical world model as a digital simulation engine embedded with comprehensive world knowledge, capable of predicting the next scene conditioned on environmental states and contextual priors. Each prediction step can be represented as a triplet (Current Scene,Navigation Mode,Prior Information), which collectively determine the evolution of the simulated environment. The model focuses on capturing the causal and spatiotemporal dynamics of the physical environment, while external inputs such as navigation modes or actions act as stimuli that perturb or impact the environment’s evolution. Hence, the world model is formulated as an interactive environment system that responds to external interventions without explicitly modeling the decision-making process that generates them. Based on this definition, we systematically analyze how video generation models have evolved towards world modeling and propose a four-generation taxonomy according to model capability, as illustrated in Figure 1.
@@ 3 Roadmap
Semantic and Navigational Interactiveness. Interactiveness is the core capability emphasized in Generation 2, termed as semantic and navigational interactiveness, marking a significant advancement in the dimensions of control flexibility, condition-video consistency, subject-centric controllability, and intrinsic model competence. Models in this generation enable flexible control, particularly conditioned on navigation modes, including actions, text commands, and predefined trajectories. They also exhibit emerging basic reasoning abilities, such as inferring motion sequences from high-level commands. In terms of condition-video consistency, Generation 2 models demonstrate notable improvements: they can generate complete and coherent motions with minimal dynamic distortion, and maintain visually appropriate and semantically aligned backgrounds. Another hallmark of this generation is basic subject-centric external control, where the model is capable of interpreting and executing control signals directed at a specific subject, such as instructing one agent to perform a sequence of actions or dynamically adjusting the viewpoint around that subject. However, these models typically operate within relatively static or simplified backgrounds. In addition, the intrinsic model capabilities have advanced significantly, as evidenced by the integration of 3D spatial understanding and stronger semantic comprehension rather than basic pixel-level understanding. Together, these advancements position Generation 2 as a transitional phase towards truly interactive and semantically grounded world modeling.
@@
Furthermore, Generation 4 achieves arbitrary spatial and temporal scale planning. In the spatial domain, the model can plan across macroscopic scales such as universe-level evolution and microscopic scales like microbial dynamics or atomic-level transitions. Similarly, in the temporal domain, the model is capable of operating across vast time scales, from long-term evolution spanning years or centuries (requiring time compression and critical event selection ability), to mid-scale physical world dynamics, down to fine-grained, high-frequency phenomena such as insect wing beats or human pupil micro-movements. This ability to plan across stochastic events and arbitrary scales represents a critical step towards building general-purpose simulation engines that align more closely with the complexity and uncertainty of the real world.
@@ 4 Methods: A Hierarchical Taxonomy
Before delving into each generation in detail, it is important to emphasize that the three core capabilities, faithfulness, interactiveness, and planning, develop in parallel rather than appearing sequentially. Each generation highlights one capability as its dominant focus, reflecting the primary research emphasis and maturity level of that stage, yet all three capabilities continue to advance simultaneously across generations. This co-evolution reflects the natural progression of video generation models towards comprehensive world modeling, where visual realism, controllability, and long-horizon reasoning are continuously refined in tandem rather than introduced one after another.
Similarly, the application domains, general scenes, robotics, autonomous driving, and gaming, analyzed in this section are not exclusive to any single generation. Instead, they overlap across all four, as the underlying modeling principles remain consistent while the level of capability and control gradually increases. We organize these applications under Generation 2 primarily because most of the recent works and open-source systems lie in this stage. Subsequent generations naturally extend these same applications towards higher-level planning, real-time interaction, and stochastic world simulation.
@@ 4.1.1 Foundation Models with Visual-Centric World Knowledge
We summarize representative pre-trained text-to-video (T2V) models [hong2022cogvideo, chen2023videocrafter1, chen2024videocrafter2, blattmann2023svd, wang2023modelscope, gen223, guo2023animatediff, zhang2024show1, wang2025lavie, qing2024higen, ma2024latte, wang2024magicvideov2, bar2024lumiere, klingai2024kling, henschel2025streamingt2v, girdhar2023emuvideo, kondratyuk2023videopoet, yuan2024instructvideo, zhou2024allegro, tian2024videotetris], along with commonly used pre-trained image-to-video (I2V) models [zhang2023i2vgen, lin2024opensoraplan, hailuo24minmax, xu2024easyanimate, klingai2024kling, blattmann2023svd, yu2023magvitv2, chen2023videocrafter1, sun2024dimensionx, kondratyuk2023videopoet, guo2024i2vadapter, team2025kling], examining their backbone architectures, video lengths, resolutions, and post-training capabilities related to interactivity and embodiment which are core abilities for serving as world models. This survey also covers both open-sourced models [ma2024latte, wang2025lavie, xu2024easyanimate, si2025repvideo, chen2024videocrafter2, guo2023animatediff, hacohen2024ltxvideo, chen2023videocrafter1, zheng2024opensora, zhang2024show1, lin2024opensoraplan, qing2024higen, wang2023modelscope, henschel2025streamingt2v, sun2024dimensionx, kondratyuk2023videopoet, zhou2024allegro] and commercial APIs [hailuo24minmax, klingai2024kling, gen223, pika23, girdhar2023emuvideo]. For instance, CogVideo [hong2022cogvideo] is widely regarded as an early open-source, large-scale pretrained T2V model that played a pivotal role in initiating the development of video foundation models. The release of Kling 1.0 [klingai2024kling], built on a DiT backbone, further advanced generative AI towards longer duration and greater visual quality.
Following the initial success of Generation 1 models in producing basic text–video consistency and plausible motion dynamics, some research shifted focus towards long and efficient video generation, sometimes at the expense of faithfulness. This led to the development of streaming video generation based on an autoregressive architecture. Specifically, LTX‑Video [hacohen2024ltxvideo] achieves semi‑real‑time generation, continuously producing video clips faster than they can be viewed, by leveraging a high‑compression Video‑VAE and a denoising transformer with full spatiotemporal attention. While notable for its generation speed, its limited fidelity of generated videos still places it within the first generation.
@@ 4.1.3 Navigation World Model
Navigation world models inviting $N$ into the input set as $I=\{T,O,Au,N\}$. In general scenes, navigation world models are primarily developed for motion control tasks, including local subject motion [burgert2025gowiththeflow, geng2025motionprompting, namekata2024sgi2v, ma2024trailblazer, jain2024peekaboo, wang2024motionctrl, wu2024draganything, qiu2024freetraj, yang2024directavideo] and camera motion [bai2025recammaster, wang2025cinemaster, bahmani2025ac3d, bahmani2024vd3d, he2024cameractrl, van2024gcd, kuang2024cvd, bai2024syncammaster]. These methods often build upon pre-trained video foundation models [blattmann2023svd, wang2023modelscope, chen2023videocrafter1, yang2024cogvideox, bar2024lumiere, guo2023animatediff], incorporating training-free control modules to enable interactive motion and camera control. However, due to current limitations in video length, generation quality, and the diversity of controllable entities, these approaches remain in the Generation 1 world models. For local motion control tasks, DragAnything [wu2024draganything] offers intuitive point‑based dragging to finely adjust subject motion, enabling precise spatial manipulation in generated videos. Peekaboo [jain2024peekaboo] elevates this to semantic‑level control, where object‑centric motion can be directed in line with textual prompts. Pushing further, Trailblazer [ma2024trailblazer] incorporates trajectory‑driven navigation, supporting complex and continuous motion generation for both subjects and cameras, advancing towards more coherent and context‑aware video navigation. In terms of camera control, existing methods [bahmani2024vd3d, bahmani2025ac3d, xu2024camco] inject camera motion trajectories, represented in Plücker coordinates, into pre-trained foundation models via ControlNet-based modules. Alternatively, methods such as GCD [van2024gcd] directly manipulate camera extrinsics, i.e., rotation and translation, as input to control camera viewpoints.